# Evaluating Time-Dependent Methods and Seasonal Effects in Code Technical Debt Prediction

### v1

@@ 3.4 Data Collection
This study included 31 projects from the previously described Technical Debt Dataset. For this study, we used all commits analyzed by SQ and extracted their commit information, SQALE index, and code smell rule violations reported by SQ.
@@ 3.4.1 Data preprocessing: Feature selection
The collected data contains 205 potential independent variables. Such a large number could have a negative impact on the models due to the so-called “curse of dimensionality”: including a large number of independent variables may drastically reduce the predictive power of a model. To identify a set of the most explicative independent variables for our dependent variable, we undertook a feature selection process that utilized well-known feature selection techniques. The aim was to categorize the potential independent variables according to their importance. These techniques are briefly described next.
@@
Variance Thresholding: A method utilizing the variance or spread of a variable in a data set for selecting model-independent variables. In a nutshell, this method calculates the variance of all the potential independent variables across the data sets and removes those with low variance as they are less informative and do not have much predictive power [26].
@@
Zero percentage method: This method eliminates predominantly empty independent variables, i.e. the percentage of emptiness is higher than a predefined threshold, and is therefore poorly informative. It is used with count data [27].
@@
Feature importance: This method determines the importance of each variable in a data set using a Random Forest (RF) algorithm. RF is an ensemble learning technique consisting of multiple decision trees, in this case, used for prediction. The measure of importance is based on the reduction in node impurities due to the splitting of the variable. Averaged over all trees, the node impurity measure of all the independent variables is calculated by the residual sum of squares. Thus, this method excludes variables with low feature importance [28].
@@
Correlation analysis: The method evaluates the degree of linear association between the independent variables and the dependent variable. Given the high skewness illustrated in Figure 1, and therefore lack of normality in our variables [29], we measure the correlational association through the non-parametric Spearman’s rank correlation coefficient [30] for each independent variable. Thus, variables with a poor informative contribution for predictive modelling are discarded [31].
@@
Understandably, one could expect that each of the selected feature selection techniques would provide different results on the initially collected issue type variables. Therefore, we ranked the independent variables in terms of their resulting importance score for each of the feature importance techniques, thus resulting in four different rankings. We used the Interquartile Range (IQR) method to select the uppermost quartile in the distribution of the rankings based on the importance scores. In a nutshell, given a set of observations, the IQR method builds underlying the distribution from the observations and selects the subset of the different quartiles in the distribution based on the goal of the analysis [32]. In our scenario, we were interested in selecting the most informative independent variables of our set of variables; therefore, we selected the quartile with the variables denoting the highest feature importance score. Additionally, for the sake of consistency in the final set of independent variables, we selected the subset of variables that showed importance in each of the four fetched quartiles. Consequently, the feature selection process resulted in a subset of 15 independent variables that were specifically selected to predict the SQALE index (dependent variable). Table 1 and Figure 1 provide a graphical and quantitative portrayal of the independent variables.
3.4.2 Data preprocessing: Transforming the raw data into time series data
@@
The time-dependent approaches used in this paper require using periodically measured time-serialized data. Given that commits are not periodical, we relied on two criteria to generate the time series. First, we focused on constructing a sufficiently long time series for each of the projects (study subjects) to provide the models with sufficient data to perform the prediction. Second, we investigated different periodicity levels for the observations in the time series. In this way, we could observe the temporal nature of the data and find the most suitable time frame between observations.
@@
Based on the described criteria we generated time series data at biweekly and monthly periodicity levels. The number of data points for each project on each periodicity level is presented in Table 2. TSA models require at least 24 observations to detect existing time-dependence factors such as the temporal trend, and consequently to be trained [23]. Therefore, increasing the time window size would reduce the number of observations for the collected projects, and require excluding projects from the study.
@@
Version control data is stochastically ordered by nature555https://www.git-scm.com/docs/git-commit, which means that some time periods (of weeks or months in our study) might not have commits despite otherwise being an active project. Consequently, serializing such data can result in missing data for some time points. Linear interpolation was used to fill in time points that lacked observations in the generated time series. This method presupposes a linear correlation between neighboring data points and calculates the values of the missing data points using the known values surrounding them [33].
@@
Consequently, the analysis was conducted using three different samples of the data: monthly time series data, biweekly time series data, and original commit data, which is chronologically ordered but not serialized in periods, and is only used with the ML models. As a potential threat to the validity of this study, we address the interpolation of artificially generated observations in Section 6. Similarly, we provide suggestions for practitioners on the remediation approach to avoid missing data in Section 4.1 to help practitioners with generating higher-quality data.
@@ 3.5 Data Analysis
In this section, we focus on the designed data analysis based on the definition of the selected models, and explain the comparison of the seasonally adjusted models against their root model counterparts, their model parameter tuning structure, and the ML models adopted. The selected time-dependent models are based on multivariate TSA modelling, a field of statistical modelling commonly used in diverse fields such as econometrics and weather forecasting for instance [34]. The multivariate nature of the models allows us to consider information about additional independent variables that can assist the model in the prediction process.
3.5.1 Setting the stage for seasonally unadjusted TSA model prediction
@@
In order to address RQ1, we examine two distinct methods to predict Code TD utilizing TSA models, ARIMAX and ARIMA+LM. As version control data is inherently non-stationary, i.e. denote a trend across time, the two methods originate from the Box-Jenkins model. better known as Auto-Regressive Integrated Moving Average (ARIMA) model. It is worth noticing that a time series of data is stationary when its predictions do not depend on factors such as trend or seasonality existing in the real data [35]. The ARIMA model is a modified version of the Autoregressive Moving Average (ARMA) model used for modelling non-stationary time series [36]. Like the ARMA model, ARIMA determines the best autoregressive parameter for the variable to be predicted (AR step) and also identifies the optimal moving average parameter (MA) [34]. In addition, ARIMA executes the integration step (I), which determines the level of differencing required for the time series of the variable in question to achieve stationarity. The methods considered are:
@@
ARIMAX: The approach used in [11] was an ARIMA model with a set of independent variables included in the model, commonly referred to as ARIMAX. This model offers a further multivariate approach to the ARIMA model by introducing additional independent variables into the model and, therefore, helping to explain the evolution of the SQALE index. This enables future SQALE index values to be explained by both its past values and those of the independent variables.
@@
ARIMA + LM: An ensemble technique previously adopted in [12] that comprises constructing a univariate ARIMA model for each independent variable in the model. Each ARIMA model considers the past values of their respective model variable as model parameters and, therefore, independently predicts the future values of each model variable. The additional step in this methodology relies on performing Linear Regression (LM) considering each of the new prediction lag values to predict the new values for the SQALE index variable.
3.5.2 Setting the stage for Seasonally Adjusted TSA model prediction
@@
In contrast with the previously presented ARIMAX and ARIMA+LM models, the seasonally adjusted models capture the impact of the seasonality pattern in the data while maintaining the model nature of the previously defined TSA model counterparts. Therefore, the analyzed seasonally adjusted methods are:
@@
SARIMAX: The Seasonally adjusted Auto-Regressive Integrated Moving Average model stands as the seasonally adjusted extension of the ARIMAX model. It addresses the seasonality of the data within the model training through the model parameter tuning [13]. Further information regarding the model parameter tuning in the TSA model is provided later in Section 3.5.3. Similarly, the SARIMAX model provides a multivariate approach compared to its univariate counterpart, SARIMA, by including independent variables in the model. This enables future SQALE index values to be explained by both its past values and those of the included independent variables.
@@
SARIMA + LM: The approach stands as the combination of the univariate Seasonally adjusted Auto-Regressive Integrated Moving Average (SARIMA) model combined with a Linear Regression model (LM). Based on the seasonality adjustment performed in the approach used in [12], a SARIMA model is built per each of the considered independent variables in the study, and thus each model independently predicts the future values of a single variable accordingly. The additional step in this methodology relies on performing Linear Regression (LM) considering each new prediction lag value of the independent variables to predict the new values for the SQALE index variable.
3.5.3 Setting the model parameter order in TSA models
@@
Similarly as in other families of predictive analysis models, for instance, Linear Models or Generalized Linear Models [29], TSA models are founded on a set of model parameters which are estimated after the model is trained. As mentioned earlier in this section, in this study we adopt as prediction models derivations from the Box-Jenkins ARIMA model [13]. Models founded on the ARIMA model follow the same model parameter structure, commonly depicted as follows:
@@
where p stands as the number of steps that a model needs to go back into the past history data to predict the next value i.e, autoregression (AR) level, q denotes the number of moving averages required (MA) and captures the short-term random fluctuations in the data, and d defines the differencing level required to achieve stationarity in the trend observed within the data and therefore be able to make predictions [37]. Models must be adjusted to these factors in order to reach stationarity on the trained data. We provide a detailed description of how parameter tuning is undergone with ARIMA-based models in the Appendix.
@@
Similarly, models capturing seasonal patterns such as SARIMA are presented with an extended set of parameters
@@
When the model is adapted to capture the seasonal patterns of the data in models such as SARIMAX and SARIMA+LM, the parameters P, D, Q are calculated as well. These capture the impact of seasonality in the model parameters explained before, which the model tunes based on the specified m cycle level in the data. In our study, we consider monthly and biweekly cycles as previously defined. Thus, the adjusted models provide an additional layer in the model parameter tuning where the seasonality component is treated.
@@
In this study, the presented model parameters are tuned when the TSA models are fitted with the training data. To determine the optimal combination of the model parameters, we conduct an iterative process in which we search for the best model parameter combination for each of the project’s time series data. Thus, for different combinations of and parameter values, we utilize the Augmented Dickey-Fuller test [38] to determine the optimal value d of the differencing model parameter. Similarly, when the seasonality effect is adjusted, the Canova-Hansen test [39] is used to determine the optimal order of seasonal differencing as well, along with different combinations of and . To execute this iterative process, we adopt the auto arima algorithm [40, 41] which defines the potential range of parameter values based on the fitted data.
@@
Thus, based on the temporal patterns of the SQALE index, the computed different model combinations of the defined parameters are assessed through Akaike Information Criterion [42] (AIC) and Bayesian Information Criterion [43] (BIC) following the state-of-the-art in statistical modelling [29]. We provide the theoretical description of the presented model parameters in Appendix Appendix: Parameter tuning of the Box-Jenkings models. Moreover, we provide the implemented logic in the shared online appendix (see Section 3.7).
3.5.4 Setting the stage for Machine Learning models

### v2

@@ 3.4 Data Collection
In this study, we were interested in studying the impact of temporal factors on the selected dependent variable, the SQALE index. As described in Section 2, code smell rule violations are directly related to the SQALE index, which is the remediation cost. Hence, we decided to collect the 205 code smell rule violations existing in the adopted data set as the initial set of independent variables for our study.
@@ 3.4.1 Data preprocessing: Feature selection
The data collected contains 205 potential independent variables. Such a large number could have a negative impact on the models due to the so-called “curse of dimensionality:” including a large number of independent variables may drastically reduce the predictive power of a model. To identify a set of the most explicative independent variables for our dependent variable, we undertook a feature selection process based on well-known feature selection techniques. The aim was to categorize the potential independent variables according to their importance. These techniques are briefly described next.
@@
Variance Thresholding: Method that utilizes the variance or spread of a variable in a data set to select model-independent variables. In summary, this method calculates the variance of all potential independent variables in the data sets. Removes those with low variance as they are less informative and do not have much predictive power [30].
@@
Zero percentage method: This method eliminates predominantly empty independent variables, i.e. the percentage of emptiness is higher than a predefined threshold and is therefore poorly informative. It is used with count data [31].
@@
Feature importance: This method determines the importance of each variable in a data set using a Random Forest (RF) algorithm. RF is an ensemble learning technique consisting of multiple decision trees, in this case, used for prediction. The measure of importance is based on the reduction in node impurities due to the splitting of the variable. Averaged across all trees, the measure of the impurity of the node of all independent variables is calculated by the residual sum of squares. Thus, this method excludes variables with low feature importance [32].
@@
Correlation analysis: The method evaluates the degree of linear association between the independent and dependent variables. Given the high skewness illustrated in Figure 1, and therefore lack of normality in our variables [33], we measure the correlational association through the non-parametric Spearman’s rank correlation coefficient [34] for each independent variable. Thus, variables with a poor informative contribution for predictive modeling are discarded [35].
@@
One could expect that each of the selected feature selection techniques would provide different results on the initial collected issue-type variables. Therefore, we ranked the independent variables in terms of their resulting importance score for each of the importance techniques of the characteristics, resulting in four different rankings. We used the Interquartile Range (IQR) method to select the upper quartile in the ranking distribution based on the importance scores. In a nutshell, given a set of observations, the IQR method builds the underlying distribution from the observations. Select the subset of the different quartiles in the distribution based on the objective of the analysis [36]. In our scenario, we were interested in selecting the most informative independent variables of our set of variables; therefore, we selected the quartile with the variables denoting the highest feature importance score. Additionally, for the sake of consistency in the final set of independent variables, we selected the subset of variables that showed importance in each of the four fetched quartiles. Consequently, the feature selection process resulted in a subset of 15 independent variables that were specifically selected to predict the SQALE index (dependent variable). Table 1 and Figure 1 provide a graphical and quantitative representation of the independent variables.
3.4.2 Data preprocessing: Transforming raw data into time series data
@@
The time-dependent approaches used in this paper require the use of periodically measured time-serialized data. Given that commits are not periodic, we relied on two criteria to generate the time series. First, we focused on constructing a sufficiently long time series for each project (study subjects) to provide the models with sufficient data to perform the prediction. Second, we investigated different periodicity levels for the observations in the time series. In this way, we could observe the temporal nature of the data and find the most suitable time frame between observations.
@@
Based on the described criteria, we generated time series data at biweekly and monthly periodicity levels. The number of data points for each project on each periodicity level is presented in Table 20. TSA models require at least 24 observations to detect existing time-dependent factors such as the temporal trend, and consequently to be trained [25]. Therefore, increasing the size of the time window would reduce the number of observations for the collected projects and would require excluding the projects from the study.
@@
Version control data are stochastically ordered by nature,555https://www.git-scm.com/docs/git-commit which means that some time periods (of weeks or months in our study) might not have commits despite being otherwise an active project. Consequently, serializing such data can result in missing data for some time points. Linear interpolation was used to fill in time points that lacked observations in the generated time series. This method presupposes a linear correlation between neighboring data points and calculates the values of the missing data points using the known values surrounding them [37].
@@
Consequently, the analysis was performed using two different samples of the data: monthly time series data and biweekly time series data. As a potential threat to the validity of this study, we address the interpolation of artificially generated observations in Section 6. Similarly, we provide suggestions for practitioners on the remediation approach to avoid missing data in Section 4.1 to help practitioners generate higher quality data.
@@
Legend: C - Closed, C* - Closed with “Other” option, L - Linkert, O - Open Ended
| RQ | Question (Q) | Type | |
| Profiling | Q1 | What is your job title or role? | O |
| Q2 | What sector does your organization belong to? | O | |
| Q3 | What is the size of your organization? | C | |
| Q4 | Where is your organization located? | C | |
| Q5 | How many years of experience do you have? | C | |
| Q6 | What types of projects do you usually work on? | C* | |
| Q7 | Do you use any agile development practices? | C | |
| Q8 | Are you familiar with the concept of technical debt? | C | |
| Q9 | If yes, what agile practice do you employ? | O | |
| Q10 | How often do you address Technical Debt-related issues? | O | |
| Q11 | What are the types of Technical Debt issues you address most frequently? | O | |
| RQ6.1 | Q12 | How much would you value a tool for Technical Debt forecasting? | L |
| Q13 | Please justify your choice in the previous question. | O | |
| Q14 | Which forecasting window combinations are acceptable and useful for your workflow? | C | |
| Q15 | Please justify your choice in the previous question. | O | |
| Q16 | Among the following options, what forecasting window combinations may you find acceptable and useful for your workflow? | C | |
| Q17 | Please justify why you chose the previous options. | O | |
| RQ6.2 | Q18 | If given the choice, would you rather choose to predict Technical Debt: Weekly, By-Weekly, Monthly, Other? | C* |
| Q19 | Please justify your choice in the previous question. | O |
3.5 Collecting the practitioners’ forecasting perceived valuable
TD is a persistent challenge in software development, affecting maintainability, quality, and long-term sustainability. Although retrospective code TD analysis is widely studied, proactive forecasting remains an emerging area with limited understanding of its perceived value in industry. To bridge this gap, we designed this survey to capture the perspectives of software professionals on the usefulness, preferred forecasting windows, and optimal scope of Code TD predictions. Our target audience consists of software engineers, architects, and technical leaders involved in technical debt management. Our objective is to evaluate whether our forecasting model is perceived as beneficial to the industry. Specifically, we want to determine the most advantageous forecasting time windows that add value to the software development lifecycle (SDLC) within an industrial setting.
3.5.1 Population Description
| Question | Response | % | |
| Q1 | Job title | Software Architect | 9 |
| Software Developer | 30 | ||
| Software Engineer | 4 | ||
| Q2 | Organization sector | Cloud Architect | 4 |
| ICT | 35 | ||
| IT Consulting | 30 | ||
| Software Engineering | 26 | ||
| Software Development | 4 | ||
| Q3 | Organization size | Large Enterprise | 65 |
| SME | 17 | ||
| Small | 13 | ||
| Start-up | 4 | ||
| Q4 | Organization location | Europe | 87 |
| North America | 4 | ||
| South America | 9 | ||
| Q5 | Years of experience | ¡ 1 | 9 |
| 1-3 | 26 | ||
| 4-7 | 30 | ||
| 8-10 | 13 | ||
| ¿ 10 | 22 | ||
| Q6 | Type of Project | AI/ML | 29 |
| Web Development | 29 | ||
| Cloud & DevOps | 12 | ||
| IoT | 8 | ||
| Embedded Systems | 8 | ||
| Mobile Development | 4 | ||
| Software Architecture & Quality | 8 |
| Question | Response | % | |
| Q7 | Use of agile practices | No | 26 |
| Yes | 74 | ||
| Q8 | Familiarity with TD | No | 13 |
| Somewhat | 17 | ||
| Yes | 70 | ||
| Q9 | Agile Practice Employed | Scrum (SCRUM, Lean + Scrum) | 56.25 |
| CI/CD, Version Control, Test Automation | 18.75 | ||
| Kanban | 6.25 | ||
| MLOps | 6.25 | ||
| Feature-Driven Development (FDD) | 6.25 | ||
| Q10 | Frequency addressing TD | Weekly | 4 |
| Per Sprint | 13 | ||
| Monthly | 13 | ||
| Quarterly | 4 | ||
| Project Start | 4 | ||
| As Needed | 9 | ||
| Occasionally | 22 | ||
| Rarely | 13 | ||
| Q11 | TD type addressed | Refactoring (general & systematic) | 26.09 |
| Testing Debt (test smells, lack of unit tests) | 17.39 | ||
| Code smells | 13.04 | ||
| Bug fixing (small, unforeseen, product bugs) | 13.04 | ||
| Architectural Debt (smells, poor planning, scalability issues) | 13.04 |
Table 3 and Table 4 present the professional experience of our the 23 interviewed experts. The surveyed population is predominantly composed of software professionals, particularly Software Developers, and is heavily concentrated in Europe. The respondents largely work in established firms within the ICT, IT Consulting, and Software Engineering sectors. They possess a range of professional experience, with a significant number employing Agile methodologies in their development practice and familiarity with code TD. Hence, our population appears to be a fair sample of the intended audience. We collected the answers anonymously and provided all participants with information regarding the GDPR act for data collection and protection. We also thoroughly followed the ACM publications policy on research involving human participants and subjects [38].
3.5.2 Questionnaire
Table 2 presents the questions used to answer RQ6. Due to space constraints, we provide the complete questionnaire in the replication package. It should be noted that, before submitting the questionnaire, we have performed a pilot questionnaire to derive the answers to the predefined closed-ended questions (C) predefined answers. Moreover, for most closed-ended questions, we provided an open-ended version (O) to account for possible missing categories or to allow for an in-depth explanation of the selection. We interviewed experts involved in our previous research and, leveraging domain knowledge, reached a consensus on the predefined answer to closed-ended questions. We did not ask the pilot’s experts to participate in the final questionnaire to avoid biases.
According to our guidelines [39, 40, 24, 41], we asked participants to answer demographic questions to obtain information about the population under examination. Thus, Q1 to Q11 refer to the personal background and professional activities of the interviewee. We asked participants to respond with predefined choices to facilitate data analysis using domain knowledge [42, 27].
For RQ6.1, we asked them for their perceived benefit in predicting the code TD. More specifically, we provide a Likert scale question (L), Q12 surveying how much would they value a tool for code TD prediction. A Likert scale is a psychometric scale commonly used in questionnaires to gauge respondents’ attitudes, opinions, or perceptions on a particular topic [43]. It typically consists of a series of statements in which respondents indicate their level of agreement or disagreement on a symmetric agree-disagree scale, usually ranging from “strongly agree” to “strongly disagree.” This method allows the quantitative measurement of people’s attitudes or feelings toward a subject. The values for each Likert scale question are available in the replication package. Furthermore, we asked the interviewee to justify the rationale for their Likert choice in Q13. In addition, we also provide two closed-ended questions, Q14 and Q16 to choose between pairs of “¡time windows, MAPE¿.” To allow the practitioner to grasp our ability to predict Code TD, we decided, based on RQ5 findings, to select a maximum time frame of 36 months, that is, 3 years. We divided such time frames into biweekly prediction time windows Q13 and monthly time window Q15 by pairing the time window with its computed MAPE (e.g., 2 weeks - 1.44 %). Since we were interested in understanding the rationale for the choice, we also asked the practitioners to further disclose their motivation for the chosen intervals in the two connected open-ended questions Q15 and Q17.
Finally, for RQ6.2, we ask whether the interviewee prefers to predict the code TD weekly or monthly Q18 based on previous comparisons of time windows and MAPE, and to justify their choice in Q19.
@@ 3.6 Data Analysis
In this section, we focus on the designed data analysis based on the definition of the selected models, and we explain the comparison of the seasonally adjusted models against their root model counterparts, their model parameter tuning structure, and the ML models adopted. The selected time-dependent models are based on multivariate TSA modeling, a field of statistical modeling commonly used in diverse fields such as econometrics and weather forecasting, for instance [44]. The multivariate nature of the models allows us to consider information about additional independent variables that can assist the model in the prediction process. Figure 2 shows the visual representation of the data analysis process performed in this study.
3.6.1 Setting the stage for seasonally unadjusted TSA model prediction (RQ1 - RQ5)
@@
We examine two distinct methods to predict code TD utilizing TSA models, ARIMAX and ARIMA+LM. As version control data is inherently non-stationary, i.e. denote a trend across time, the two methods originate from the Box-Jenkins model. better known as Auto-Regressive Integrated Moving Average (ARIMA) model. It is worth noticing that a time series of data is stationary when its predictions do not depend on factors such as trend or seasonality existing in the real data [45]. The ARIMA model is a modified version of the Autoregressive Moving Average (ARMA) model used for modeling non-stationary time series [46]. Like the ARMA model, ARIMA determines the best autoregressive parameter for the variable to be predicted (AR step) and also identifies the optimal moving average parameter (MA) [44]. In addition, ARIMA executes the integration step (I), which determines the level of differencing required for the time series of the variable in question to achieve stationarity. The methods considered are the following.
@@
ARIMAX: The approach used in [11] was an ARIMA model with a set of independent variables included in the model, commonly referred to as ARIMAX. This model offers a further multivariate approach to the ARIMA model by introducing additional independent variables into the model and, therefore, helping to explain the evolution of the SQALE index. This enables future SQALE index values to be explained both by its past values and by those of the independent variables.
@@
ARIMA + LM: An ensemble technique previously adopted in [12] that comprises constructing a univariate ARIMA model for each independent variable in the model. Each ARIMA model considers the past values of their respective model variable as model parameters and therefore independently predicts the future values of each model variable. The additional step in this methodology relies on the performance of linear regression (LM) considering each of the new prediction lag values to predict the new values for the SQALE index variable.
3.6.2 Setting the stage for Seasonally Adjusted TSA model prediction (RQ3,4)
@@
In contrast with the previously presented ARIMAX and ARIMA+LM models, the seasonally adjusted models capture the impact of the seasonality pattern in the data while maintaining the model nature of the previously defined TSA model counterparts. Therefore, the analyzed seasonally adjusted methods are as follows.
@@
SARIMAX: The Seasonally adjusted Auto-Regressive Integrated Moving Average model stands as the seasonally adjusted extension of the ARIMAX model. Addresses the seasonality of the data within model training through model parameter tuning [13]. More information on the adjustment of the model parameters in the TSA model is provided in Section 3.6.3. Similarly, the SARIMAX model provides a multivariate approach compared to its univariate counterpart, SARIMA, by including independent variables in the model. This enables future SQALE index values to be explained both by its past values and by those of the included independent variables.
@@
SARIMA + LM: The approach is characterized by the combination of the univariate seasonally adjusted integrated moving average (SARIMA) model combined with a linear regression model (LM). Based on the seasonality adjustment performed in the approach used in [12], a SARIMA model is built for each of the independent variables considered in the study, and therefore each model independently predicts future values of a single variable accordingly. The additional step in this methodology relies on the implementation of linear regression (LM) considering each new prediction lag value of the independent variables to predict the new values for the SQALE index variable.
3.6.3 Setting the model parameter order in TSA models (RQ1 - RQ5)
@@
Similarly to other families of predictive analysis models, for instance, Linear Models or Generalized Linear Models [33], TSA models are founded on a set of model parameters that are estimated after the model is trained. As mentioned earlier, in this study we adopt as prediction models derivations from the Box-Jenkins ARIMA model [13]. Models founded on the ARIMA model follow the same model parameter structure, commonly depicted as
@@
where p stands for the number of steps that a model needs to go back to the history data to predict the next value, that is, the level of autoregression (AR), q denotes the number of moving averages required (MA) and captures the short-term random fluctuations in the data, and d defines the differencing level required to achieve stationarity in the trend observed within the data and therefore be able to make predictions [47]. Models must be adjusted to these factors to reach stationarity in trained data. We provide a detailed description of how parameter tuning is undergone with ARIMA-based models in the appendix.
@@
Similarly, models that capture seasonal patterns, such as SARIMA, are presented with an extended set of parameters
@@
When the model is adapted to capture the seasonal patterns of the data in models such as SARIMAX and SARIMA+LM, the parameters P, D, Q are also calculated. These capture the impact of seasonality on the model parameters explained above, which the model adjusts based on the specified m cycle level in the data. In our study, we consider monthly and biweekly cycles as previously defined. Thus, the adjusted models provide an additional layer in the model parameter tuning where the seasonality component is treated.
@@
In this study, the model parameters presented are tuned when the TSA models are fitted with the training data. To determine the optimal combination of the model parameters, we conducted an iterative process in which we searched for the best combination of model parameters for each of the project’s time series data. Thus, for different combinations of the values of the parameters and , we use the augmented Dickey-Fuller test [48] to determine the optimal value d of the diffencing model parameter. Similarly, when the seasonality effect is adjusted, the Canova-Hansen test [49] is used to determine the optimal order of seasonal differentiation , along with different combinations of and . To execute this iterative process, we adopt the Auto-arima algorithm [50, 51] which defines the potential range of parameter values based on the fitted data. We provide a detailed description of the algorithm functioning in the appendix.
@@
Based on the temporal patterns of the SQALE index, the different combinations of models computed for the defined parameters are evaluated using the Akaike Information Criterion [52] (AIC) and the Bayesian Information Criterion [53] (BIC) following the latest in statistical modelling [33]. We provide a theoretical description of the model parameters presented in the appendix. Moreover, we provide the logic implemented in the shared online package (see Section 9).
3.6.4 Setting the stage for Machine Learning models (RQ2 - RQ4)